# IMRL: Integrating Visual, Physical, Temporal, and Geometric Representations for Enhanced Food Acquisition

### v1

@@ Abstract
Robotic assistive feeding holds significant promise for improving the quality of life for individuals with eating disabilities. However, acquiring diverse food items under varying conditions and generalizing to unseen food presents unique challenges. Existing methods that rely on surface-level geometric information (e.g., bounding box and pose) derived from visual cues (e.g., color, shape, and texture) often lacks adaptability and robustness, especially when foods share similar physical properties but differ in visual appearance. We employ imitation learning (IL) to learn a policy for food acquisition. Existing methods employ IL or Reinforcement Learning (RL) to learn a policy based on off-the-shelf image encoders such as ResNet-50. However, such representations are not robust and struggle to generalize across diverse acquisition scenarios. To address these limitations, we propose a novel approach, IMRL (Integrated Multi-Dimensional Representation Learning), which integrates visual, physical, temporal, and geometric representations to enhance the robustness and generalizability of IL for food acquisition. Our approach captures food types and physical properties (e.g., solid, semi-solid, granular, liquid, and mixture), models temporal dynamics of acquisition actions, and introduces geometric information to determine optimal scooping points and assess bowl fullness. IMRL enables IL to adaptively adjust scooping strategies based on context, improving the robot’s capability to handle diverse food acquisition scenarios. Experiments on a real robot demonstrate our approach’s robustness and adaptability across various foods and bowl configurations, including zero-shot generalization to unseen settings. Our approach achieves improvement up to $35\%$ in success rate compared with the best-performing baseline. More details can be found on our website https://ruiiu.github.io/imrl/.
@@ I Introduction
While there has been progress in food acquisition for robotic assistive feeding, several limitations persist. Some existing systems [1, 2, 3, 4] rely on predefined, heuristic-based, or hard-coded primitive actions that lack generalization across different food types and configurations. Other works [5, 6, 7, 12, 8, 13, 14] incorporate visual perception to handle certain aspects of food acquisition; however, these methods mainly focus on learning information like bounding boxes and food poses from visual cues. They do not capture the food’s physical properties, which are crucial for effective manipulation. Additionally, the acquisition of semi-solid and liquid foods remains relatively underexplored in assistive feeding, with most research focusing on solid or granular foods [5, 15, 8].
We use imitation learning (IL) and select Behavior Cloning (BC) as the IL algorithm to learn a policy for food acquisition. BC has been employed in previous works [15, 13, 14] for similar tasks. However, standard BC often lacks robustness and struggles to generalize across diverse acquisition scenarios. The effectiveness of BC is heavily dependent on the quality of the representations used. Although we only focus on BC, since our advancement is in the representations, we expect that this will be just as useful in other downstream approaches such as RL [16, 17].
Figure 1: Comparison of standard BC and our approach for food acquisition. The standard BC (top) processes robot observations through an off-the-shelf encoder (e.g., ResNet-50), generating actions that are not robust and generalizable. In contrast, our proposed approach IMRL (bottom) utilizes visual, physical, temporal andd geometric representation learning to develop richer and more informative representations to enhance the robustness and generalizability of BC for food acquisition.
To address these gaps, our approach focuses on enhancing the robustness and generalizability of BC by developing richer and more informative representations for food acquisition in assistive feeding. We propose a novel approach that integrates visual, physical, temporal, and geometric representations to achieve that. Our pipeline, depicted in Figure 1, incorporates visual and physical representation to capture food types and physical properties. The core idea is to develop representations that not only distinguish between different food types but also capture their physical properties. This capability is crucial for downstream BC tasks, as foods with similar physical properties but different shapes or colors may necessitate similar scooping strategies (e.g., water and milk), while foods with the same color but different physical properties may require distinct strategies (e.g., milk and rice).
@@
To summarize, the key contributions of our work:
  • We introduce a novel approach that integrates visual, physical, temporal, and geometric representations to provide a richer understanding of foods. This advanced representation captures details that are critical for effective food acquistion, going beyond surface-level visual information.
@@ III-B Representation Learning
Figure 3: Visual and physical representation learning. Given a food image dataset, we not only learn visual semantics of different food types with CE loss, but also food physical properties with Triplet loss.
@@ III-B2 Physical Properties
We apply contrastive learning with triplet loss to ensure that representations of food items with similar physical properties are pulled closer together, while those with differing properties are pushed further apart.
We define the triplet as following: $\mathcal{A}$ (anchor), a sample whose representation that we want to anchor (e.g., milk); $\mathcal{P}$ (positive), a sample that is similar to the anchor in terms of the physical property (e.g., water, which shares the liquid property with milk); $\mathcal{N}$ (negative), a sample that is different from the anchor (e.g., rice, which is granular).
We use the L2 norm to measure the dissimilarity between feature representations: $d(f(x_{i}),f(x_{j}))=||f(x_{i})-f(x_{j})||_{2}$, where $f$ is the feature extractor that maps an input image $x$ to its feature representation $f(x)$.
Mathematically, we define the triplet loss for a triplet $(\mathcal{A},\mathcal{P},\mathcal{N})$ as: $\mathcal{L}(\mathcal{A},\mathcal{P},\mathcal{N})=\max\big{(}0,d(f(\mathcal{A}),f(\mathcal{P}))-d(f(\mathcal{A}),f(\mathcal{N}))+\alpha\big{)}$, where $\alpha$ is the margin, a hyperparameter that defines how much closer the positive should be compared to the negative.
The objective is to make the distance between the anchor and positive $d(f(\mathcal{A}),f(\mathcal{P}))$ smaller than the distance between the anchor and negative $d(f(\mathcal{A}),f(\mathcal{N}))$ by at least a margin $\alpha$. The overall triplet loss is: $\mathcal{L}_{tri}=\mathbb{E}_{D}[\mathcal{L}(\mathcal{A},\mathcal{P},\mathcal{N})]$.
@@ III-B4 Geometric Information
We first use SAM [28] to segment the desired food item and generate a binary mask over the food. SAM is known for its high-quality segmentation and adaptability without retraining.
We then compute the expectation over the mask to determine the centroid position of the food, which serves as a reference point to initialize the search for the optimal scooping point.
To find the optimal scooping point, we calculate local density for each point $(x_{i},y_{i})$ inside the food mask $(M(x_{i},y_{i})=1)$, we define a local neighborhood $N_{r}(x_{i},y_{i})$ of size $r\times r$ centered around $(x_{i},y_{i})$. The local density $\rho(x_{i},y_{i})$ for that point is given by: $\rho(x_{i},y_{i})=\frac{\sum_{(x_{j},y_{j})\in N_{r}(x_{i},y_{i})}M(x_{j},y_{j})}{|N_{r}(x_{i},y_{i})|}$, where $|N_{r}(x_{i},y_{i})|$ is the total number of pixels in the neighborhood $N_{r}(x_{i},y_{i})$. This equation computes the propertion of food pixels in the local region, effectively providing a meaure of food density around point $(x_{i},y_{i})$.
The optimal scooping point $(x^{*},y^{*})$ is the point inside the mask that maximizes the local density: $(x^{*},y^{*})=\arg\max_{(x_{i},y_{i})}\{\rho(x_{i},y_{i})|M(x_{i},y_{i})=1\}$.
@@ III-C Behavior Cloning
As mentioned earlier, the primary objective of this paper is to learn a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that guides a robot’s actions for food acquisition with BC.
The BC loss is a CE loss and is expressed as following: $\mathcal{L}=-\mathbb{E}_{\mathcal{D}}\big{[}\log\pi_{\theta}(a_{t}|s_{t-k:t})\big{]}$, where $\theta$ is the parameters of the policy, $a_{t}$ is the expert action. Minimizing the BC loss $\mathcal{L}$ allows us to optimize the scooping actions for effective food manipulation.
Unlike naive BC, we develop a richer and more informative multi-dimensional representation $z$ to enhance the robustness and generalizability of BC. This representation $z$ ensures that the learned policy can tailor scooping strategies appropriately based on different contexts.
@@ IV-A Experimental Setup
To answer question (2), we perform zero-shot generalization testing on unseen bowls and food variants. We evaluate the performance of IMRL in these new contexts and compare it against the baseline methods.
To address questions (3) and (4), we conduct ablation studies to validate our model’s design and assess the effectiveness of visual, physical, temporal, and geometric representations. These studies help determine how each module contributes to overall performance. Additionally, we perform a qualitative analysis to evaluate the advantages of our learned visual and physical representations.
@@ IV-B Data Collection
We collect a set of demonstrations as training data. Our dataset consists of $N=30$ robot trajectories obtained through teleoperation using Gello [30]. Each trajectory $i$ includes pairs of RGB images and robot joint angles recorded throughout the scooping process. Specifically, we used a white circular bowl with volume 0.74 qt containing three types of food: granular cereals, semi-solid jello, and liquid water. Notably, we did not collect data for other bowls of different sizes, shapes and colors or food variants.
@@ IV-C Baselines
We implemented several baselines for comparison: Vis: we used the method of Sundaresan et al. [12] with only visual information, as our setup does not include haptic data. SCONE [15]: we applied this method without depth information, as our setup lacks depth images. Pretrained Feature Extractors without Representation Learning: we evaluated multiple methods using pretrained models to extract features, including: ImageNet pretrained ResNet-50 [23], DINO-v2 [19], MoCo-v3 [21].
@@ IV-E Generalization Testing
We conduct zero-shot generalization testing on unseen scenarios to answer question (2). These scenarios involve various bowl types, including a large blue circular bowl, a small blue circular bowl, and a transparent square bowl, as well as different food items such as granular rice, black beans, green beans, yellow beans, and milk. To further assess the generalization capability of IMRL in handling foods with similar physical properties despite their differing color appearances, we mix black, green, and yellow beans and compare the success rate of IMRL with other baselines for acquiring unseen food. As shown in Fig. 5(d), IMRL performs comparably well on unseen foods, while the performance of other baselines decreases significantly, highlighting the limitation of standard BC methods in terms of robustness and generalizability.
@@ V Conclusions
However, there are some limitations to our approach. Our framework relies solely on visual cues, limiting performance in scenarios where haptic or tactile feedback could offer further precision in handling food. Future work could focus on incorporating multimodal sensory inputs, such as force feedback or depth sensors, to capture more nuanced physical interactions between the robot and food items.

### v2

@@ Abstract
Robotic assistive feeding holds significant promise for improving the quality of life for individuals with eating disabilities. However, acquiring diverse food items under varying conditions and generalizing to unseen food presents unique challenges. Existing methods that rely on surface-level geometric information (e.g., bounding box and pose) derived from visual cues (e.g., color, shape, and texture) often lacks adaptability and robustness, especially when foods share similar physical properties but differ in visual appearance. We employ imitation learning (IL) to learn a policy for food acquisition. Existing methods employ IL or Reinforcement Learning (RL) to learn a policy based on off-the-shelf image encoders such as ResNet-50. However, such representations are not robust and struggle to generalize across diverse acquisition scenarios. To address these limitations, we propose a novel approach, IMRL (Integrated Multi-Dimensional Representation Learning), which integrates visual, physical, temporal, and geometric representations to enhance the robustness and generalizability of IL for food acquisition. Our approach captures food types and physical properties (e.g., solid, semi-solid, granular, liquid, and mixture), models temporal dynamics of acquisition actions, and introduces geometric information to determine optimal scooping points and assess bowl fullness. IMRL enables IL to adaptively adjust scooping strategies based on context, improving the robot’s capability to handle diverse food acquisition scenarios. Experiments on a real robot demonstrate our approach’s robustness and adaptability across various foods and bowl configurations, including zero-shot generalization to unseen settings. Our approach achieves an improvement up to $35\%$ in success rate compared with the best-performing baseline. More details can be found on our website https://ruiiu.github.io/imrl.
Figure 1: Comparison of standard BC and our approach for food acquisition. The standard BC (top) processes robot observations through an off-the-shelf encoder (e.g., ResNet-50), generating actions that are not robust and generalizable. In contrast, our proposed approach IMRL (bottom) utilizes visual, physical, temporal and geometric representation learning to develop richer and more informative representations to enhance the robustness and generalizability of BC for food acquisition.
@@ I Introduction
Despite progress in food acquisition for robotic assistive feeding, several limitations persist. Some systems [1, 2, 3, 4] rely on predefined, heuristic-based, or hard-coded primitive actions that lack generalization across different food types and configurations. Other works [5, 6, 7, 12, 8, 13, 14] integrate visual perception to handle food acquisition but primarily focus on extracting information such as bounding boxes and food poses from visual cues. Capturing the food’s physical properties, crucial for effective manipulation, often requires haptic sensors, as seen in [7, 12]. Additionally, the acquisition of semi-solid and liquid foods remains relatively underexplored in assistive feeding, with most research focusing on solid or granular foods [5, 15, 8].
We employ imitation learning (IL) and use Behavior Cloning (BC) as the IL algorithm to learn a policy for food acquisition. BC has been employed in previous works [15, 13, 14] for similar tasks. However, standard BC often lacks robustness and struggles to generalize across diverse acquisition scenarios. The effectiveness of BC is heavily dependent on the quality of the representations used. Although we only focus on BC, since our advancement is in the representations, we expect that this will be just as useful in other downstream approaches such as RL [16, 17].
To address these gaps, our approach focuses on enhancing the robustness and generalizability of BC by developing richer and more informative representations for food acquisition in assistive feeding. We propose a novel approach that integrates visual, physical, temporal, and geometric representations to achieve that. As illustrated in Fig. 1, our pipeline incorporates visual and physical representations to capture both food types and their physical properties. The core idea is to develop representations that not only distinguish between different food types but also encode their physical characteristics. This is crucial for downstream BC tasks, as foods with similar physical properties but different appearances (e.g., water and milk) may necessitate similar scooping strategies, while foods with the same color but different physical properties (e.g., milk and rice) may require distinct strategies. Unlike prior works [7, 12] that rely on explicit haptic sensors to collect physical data, our approach learns an a priori representation to infer food physical properties just from visual data.
@@
To summarize, the key contributions of our work are as follows:
  • We introduce a novel approach IMRL that integrates visual, physical, temporal, and geometric representations to provide a richer understanding of foods. This advanced representation captures details that are critical for effective food acquisition, going beyond surface-level visual information.
@@ III-B Representation Learning
Figure 3: Visual and physical representation learning. Given a food image dataset, we not only learn visual semantics of different food types with a cross-entropy loss, but also food physical properties with a triplet loss.
@@ III-B2 Physical Properties
We apply contrastive learning with triplet loss to ensure that representations of food items with similar physical properties are pulled closer together, while those with differing properties are pushed further apart. We define a triplet as following: $\mathcal{A}$ (anchor), a sample whose representation that we want to anchor (e.g., milk); $\mathcal{P}$ (positive), a sample that is similar to the anchor in terms of the physical property (e.g., water, which shares the liquid property with milk); $\mathcal{N}$ (negative), a sample that is different from the anchor (e.g., rice, which is granular). We use the L2 norm to measure the dissimilarity between feature representations: $d(f(x_{i}),f(x_{j}))=||f(x_{i})-f(x_{j})||_{2}$, where $f$ is the feature extractor that maps an input image $x$ to its feature representation $f(x)$. Mathematically, we define the triplet loss for a triplet $(\mathcal{A},\mathcal{P},\mathcal{N})$ as: $\mathcal{L}(\mathcal{A},\mathcal{P},\mathcal{N})=\max\big{(}0,d(f(\mathcal{A}),f(\mathcal{P}))-d(f(\mathcal{A}),f(\mathcal{N}))+\alpha\big{)}$, where $\alpha$ is the margin, a hyperparameter that defines how much closer the positive should be compared to the negative. The objective is to make the distance between the anchor and positive $d(f(\mathcal{A}),f(\mathcal{P}))$ smaller than the distance between the anchor and negative $d(f(\mathcal{A}),f(\mathcal{N}))$ by at least a margin $\alpha$. The overall triplet loss is: $\mathcal{L}_{tri}=\mathbb{E}_{D}[\mathcal{L}(\mathcal{A},\mathcal{P},\mathcal{N})]$.
@@ III-B4 Geometric Information
We first use SAM [28] to segment the desired food item and generate a binary mask over the food. SAM is known for its high-quality segmentation and adaptability without retraining. We then compute the expectation over the mask to determine the centroid position of the food, which serves as a reference point to initialize the search for the optimal scooping point. To find the optimal scooping point, we calculate local density for each point $(x_{i},y_{i})$ inside the food mask $(M(x_{i},y_{i})=1)$, we define a local neighborhood $N_{r}(x_{i},y_{i})$ of size $r\times r$ centered around $(x_{i},y_{i})$. The local density $\rho(x_{i},y_{i})$ for that point is given by: $\rho(x_{i},y_{i})=\frac{\sum_{(x_{j},y_{j})\in N_{r}(x_{i},y_{i})}M(x_{j},y_{j})}{|N_{r}(x_{i},y_{i})|}$, where $|N_{r}(x_{i},y_{i})|$ is the total number of pixels in the neighborhood $N_{r}(x_{i},y_{i})$. This equation computes the proportion of food pixels in the local region, effectively providing a measure of food density around point $(x_{i},y_{i})$. The optimal scooping point $(x^{*},y^{*})$ is the point inside the mask that maximizes the local density: $(x^{*},y^{*})=\arg\max_{(x_{i},y_{i})}\{\rho(x_{i},y_{i})|M(x_{i},y_{i})=1\}$.
@@ III-C Behavior Cloning
As mentioned earlier, the primary objective of this paper is to learn a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that guides a robot’s actions for food acquisition with BC. The BC loss is a Negative Log-Likelihood loss and is expressed as following: $\mathcal{L}=-\mathbb{E}_{\mathcal{D}}\big{[}\log\pi_{\theta}(a_{t}|s_{t-k:t})\big{]}$, where $\theta$ is the parameters of the policy, $a_{t}$ is the expert action. Minimizing the BC loss $\mathcal{L}$ allows us to optimize the scooping actions for effective food manipulation. Unlike naive BC, we develop a richer and more informative multi-dimensional representation $z$ to enhance the robustness and generalizability of BC. We finetune $z$ during the training of BC with robot data. This representation enables the learned policy to tailor scooping strategies appropriately based on different contexts.
@@ IV-A Experimental Setup
To answer question (2), we perform zero-shot generalization testing on unseen bowls and food variants not encountered during robot demonstrations. We evaluate the performance of IMRL in these new contexts and compare it against the baseline methods. To address questions (3) and (4), we conduct ablation studies to validate our model’s design and assess the effectiveness of visual, physical, temporal, and geometric representations. These studies help determine how each module contributes to overall performance. Additionally, we perform a qualitative analysis to evaluate the advantages of our learned visual and physical representations.
@@ IV-B Data Collection
We first collect a diverse set of labeled food images from the internet for representation learning. We then collect a set of robot demonstrations as training data for BC. The BC dataset consists of $N=30$ robot trajectories obtained through teleoperation using Gello [30]. Each trajectory includes pairs of RGB images and robot joint angles recorded throughout the scooping process. Specifically, we used a white circular bowl with volume 0.74 qt containing three types of food: granular cereals, semi-solid jello, and liquid water. Notably, we did not collect data for other bowls of different sizes, shapes and colors or food variants.
@@ IV-C Baselines
We implemented several baselines for comparison: Vis: we used the method of Sundaresan et al. [12] with only visual information, as our setup does not include haptic data. SCONE [15]: we applied this method with RGB images. Pretrained Feature Extractors without Representation Learning: we evaluated multiple methods using pretrained models to extract features, including: ImageNet pretrained ResNet-50 [23], DINO-v2 [19], MoCo-v3 [21].
@@ IV-E Generalization Testing
We conduct zero-shot generalization testing on unseen scenarios from robot demonstrations to answer question (2). These scenarios involve various bowl types, including a large blue circular bowl, a small blue circular bowl, and a transparent square bowl, as well as different food items such as granular rice, black beans, green beans, yellow beans, and milk. To further assess the generalization capability of IMRL in handling foods with similar physical properties despite their differing color appearances, we mix black, green, and yellow beans and compare the success rate of IMRL with other baselines for acquiring unseen food. As shown in Fig. 5(d), IMRL performs comparably well on unseen foods, while the performance of other baselines decreases significantly, highlighting the limitation of standard BC methods in terms of robustness and generalizability.
@@ V Conclusions
Limitations and Future Work. Despite the advantages of IMRL, there are some limitations. We evaluate the quantity of food scooped after a certain number of scooping attempts; however, maximizing the amount of food scooped is not necessarily ideal, as individual preferences vary. Future work could explore conditioning the learned policy on a desired amount of food. Additionally, incorporating multimodal sensory inputs, such as force feedback or depth sensors, could capture more nuanced physical interactions between the robot and food items.